# Non-iterative triples for transcorrelated coupled cluster

### v1

@@ Abstract
We present an implementation of a perturbative triples correction for coupled cluster including single and double excitations based on the transcorrelated Hamiltonian. Transcorrelation introduces explicit electron correlation in the electronic Hamiltonian through similarity transformation with a correlation factor. Due to this transformation, the transcorrelated Hamiltonian includes up to three-body couplings and becomes non-hermitian. Since conventional coupled cluster is solved by projection, it is well suited to harbor non-hermitian Hamiltonians. The arising three-body operator, however, creates a huge memory bottleneck and increases the runtime scaling of the coupled cluster equations. As it has been shown that the three-body operator can be approximated, by expressing the Hamiltonian in the normal-ordered form, we investigate this approximation for the perturbative triples correction. Results are compared with code-generation based transcorrelated coupled cluster up to quadruple excitations.
@@ 1 Introduction
The accurate description of electron correlation remains one of the greatest challenges in quantum chemistry. Within a one-particle picture, the energy contribution of electron correlation can be defined as the difference between the exact full configuration interaction (FCI), and the mean-field Hartree–Fock (HF) energy. One of the most powerful approaches to recover the correlation energy is the coupled cluster1, 2, 3, 4 (CC) approach. In CC, the wave function is parameterized by an exponential cluster operator acting on a reference wave function (usually the HF determinant). In practice, the respective wave function expansion is truncated after single and double excitations5, 6 (CCSD) to avoid the exponential scaling. The inclusion of full triple excitations (CCSDT)7, 8 can significantly improve on the accuracy of the correlation energy9, but accounting for it is computationally too demanding for most system sizes. Hence, triples are usually approximated by perturbative approaches10 (in particular in the CCSD(T) model) as a good compromise between cost and accuracy.
@@
The size of the one-particle basis, given by a set of Gaussian functions, determines the overall quality of the a calculation. With an increasing number of Gaussians, the basis set eventually converges to the complete basis set limit. However, due to the high computational scaling of most ab initio methods, it is not always feasible to apply large basis sets.
Because it enforces the electron-electron cusp 11, 12, the explicit inclusion of the electronic distance in the wave function can accelerate the basis set convergence11, therefore increasing the accuracy of calculations with a given basis set. For most applications, the Ansätze of the / family13, 14, 15 provide a way to explicitly correlate electrons within the wave function. However, a major drawback of these methods is the non-trivial generalization to higher than doubles excitations. In most implementations, the perturbative triples correction to CCSD- is evaluated by the standard CCSD(T) formula (see Eq. (25) below), which leads to the finite basis set error dominating other potential errors. Werner and coworkers addressed this issue, by scaling the perturbative triples correction by the basis set correction of Møller-Plesset pertubation theory, which assumes the same convergence rate to the complete basis set limit and breaks the size-consistency16, 17. The size-consistency issue was later addressed by Kallay et al.18. A different approach exists, which introduced an explicitly correlated component into the expression for the triples correction19, 20. However, this approach did not yield improvements comparable to CCSD-.
A different approach to explicitly correlate electrons is transcorrelation, introduced by Boys and Handy21, which we adopt in this study. In contrast to including the inter-electronic distance into the wave function, the Hamiltonian is similarity transformed by a correlation factor. As a result, transcorrelation is generally applicable, even to multi-configurational wave functions 22, 23, 24, 25. However, the resulting transcorrelated Hamiltonian is no longer hermitian and includes a three-body operator. Recently, the transcorrelation idea has experienced a vivid revival26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 24, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 25, 59, 60, 61, 62, 63.
In the context of CC, the non-hermiticity does not impose any issues, since the Hamiltonian is similarity transformed by the exponential cluster operator, resulting already in a non-hermitian Hamiltonian. Nevertheless, the full treatment of the three-body operator increases the computational cost and memory scaling. Ten-no combined the transcorrelated Hamiltonian with linearized CC64, where many contractions involving the three-body operator were neglected. Later Alavi and coworkers applied CCSD to the transcorrelated Hamiltonian (TC-CCSD) for the homogeneous electron gas49, atoms50 and molecules51. In order to deal with the three-body couplings, they introduced an approximation based on the normal-ordered Hamiltonian51. In this approximation, the full TC Hamiltonian is normal-ordered with respect to a single reference determinant and subsequently the three-body integrals are neglected. Within this procedure, mean-field contributions are transferred from the three-body operator to two-, one-, and zero-body operators. In the same paper, they introduced a similar approximation based on the -dressed Hamiltonian65, 66, which, however, still suffers from the huge memory scaling. In the -dressed Hamiltonian approximation, all remaining three-body terms are neglected, meaning that effective intermediates are pre-contracted with the three-body operator before the operator is neglected. Another approach67 to reduce the drawbacks of the three-body operator is based on the generalized normal-ordering68, 69 of the three-body operator, which reduces the scaling to .
For higher-order CC models, the explicit treatment of the three-body operator results in numerous possible diagrams. A derivation of the CCSD equation for Hamiltonians including three-body couplings was given by Piecuch and coworkes in the context of nuclear structure theory70. The total number of CCSD diagrams for a Hamiltonian including three-body couplings increases from 48 to 116. Hence, tools for automatic code contractions become important for the implementation of TC-CC with higher than doubles excitations.
@@
In this work, we derive a non-iterative triples approximation for TC-CCSD(T) including three-body operators and the normal-order approximation. A first CCSD(T) implementation for Hamiltonians including a three-body operator71 was derived based on the CCSD(T)72, 73, 74 method. Later, Kats et al. derived the CCSD(T) method for the transcorrelated Hamiltonian62, however without the inclusion of the full three-body operator.
We implemented TC-CCSDT and TC-CCSDTQ for the full transcorrelated Hamiltonian with the Wick&d75, 76 library to generate the required CC tensor contractions. In the following section, we first briefly review the conventional non-iterative triples approximation and the transcorrelated Hamiltonian. Afterwards, we show the required diagrams including three-body couplings for perturbative triples. The correlation factor applied in this work is introduced in section 3. We compare these TC-CC models for Be, LiH and Be2 in section 4.
@@ 2 Theory
If not stated otherwise, we assume spin-orbital labels for all operators. For convenience, we summarize our notation in Table 1.
| Symbol | Meaning |
| occupied spin-orbital indices | |
| virtual spin-orbital indices | |
| general spin-orbital indices | |
| denotes a normal-ordered string of second quantized operators | |
| electron repulsion integrals in physics notation | |
| anti-symmetrized two-body integrals, e.g. |
@@
is the mean-field energy contribution of the reference determinant, including the nuclear repulsion energy . is an element of the Fock matrix
where comprises all one-body operators, i.e. the kinetic energy and electron-nuclear interactions. In order to distinguish from the transcorrelated Hamiltonian, the former quantity is referred to as the conventional Hamiltonian. We consistently map the indices of the creation and annihilation operators by placing them in the superscript and subscript, respectively, in the corresponding tensor (bra and ket of anti-symmetrized integrals, respectively).
We employ the usual diagrammatic representation77, 78, 79, where dashed lines represent operators. For example, two outgoing arcs connected to an “x” correspond to a one-body operator, two outgoing arcs connected to another set of two outgoing arcs represent a two-body operator. Solid horizontal lines represent amplitudes, where the number of vertices corresponds to the excitation degree of the amplitude. The dotted line represents the denominator constructed from the diagonal entries of the Fock operator; for example
is the three-body denominator. The coupled cluster amplitudes are denoted by the usual symbols , , , and so forth. For convenience, we will employ both an algebraic and diagramatic representation in this work.
@@ 2.1 Conventional perturbative triples
In this section, we re-derive the conventional perturbative triples correction to the energy in the formalism of Stanton80. In subsequent section, this framework is extended to incorporate the three-body operators in the electronic Hamiltonian.
As described by Stanton in Ref. 80, CCSD can be transformed into an eigenvalue problem of ,
where the eigenvector is the reference determinant . However, since the similarity transformation is not unitary, is not hermitian. Hence, the left eigenvector is not given by
In order to evaluate the left eigenvector explicitly, the so-called lambda equations have to be solved, so that
with where is a de-excitation operator. Now, instead of , is the zeroth-order Hamiltonian with and as left- and right-hand states, respectively. The perturbative expansion of the similarity transformed Hamiltonian according to Stanton80 results in the leading-order correction to the energy
where S, D, and T correspond to singly, doubly, and triply excited determinants, respectively, and D3 is a diagonal matrix with the inverse of the elements defined in Eq. (4) on the diagonal. This perturbative expansion leads to two different sets of amplitudes which can be evaluated according to
@@
Note that due to the commutator in Eq. (11) only connected diagrams can be created, while in Eq. (12) also disconnected diagrams appear. The tensor contractions for the first and second intermediate have the following form
@@
respectively. Note that the last term in the latter equation vanishes due to the Brillouin condition in the case of conventional Hamiltonian, but this is no longer true for the transcorrelated Hamiltonian considered in the subsequent sections. Because of the anti-symmetry of the two-body integrals and the approximation of Eqs. (20) and (24) are equivalent. Therefore, in practice, the latter two diagrams in Eq. (20) are added to Eq. (15), so that the first tensor contraction must be evaluated only once. The energy contribution of the non-iterative triples correction for the conventional Hamiltonian can be written as
The skeleton diagrams in Eq. (29) can be derived by contracting the two intermediates with each other.
@@ 2.2 Transcorrelated Hamiltonian
The TC Hamiltonian can be derived by similarity transformation of the conventional Hamiltonian with a Jastrow factor81 of the form
resulting in the additional two-body term and the three-body term . For the sake of completeness, we summarize some of the important aspects of the TC Hamiltonian. In the following, we redefine

### v2

@@ Abstract
We present an implementation of a perturbative triples correction for the coupled cluster ansatz including single and double excitations based on the transcorrelated Hamiltonian.
Transcorrelation introduces explicit electron correlation in the electronic Hamiltonian through similarity transformation with a correlation factor.
Due to this transformation, the transcorrelated Hamiltonian includes up to three-body couplings and becomes non-Hermitian.
Since the conventional coupled cluster equations are solved by projection, it is well suited to harbor non-Hermitian Hamiltonians.
The arising three-body operator, however, creates a huge memory bottleneck and increases the runtime scaling of the coupled cluster equations.
As it has been shown that the three-body operator can be approximated, by expressing the Hamiltonian in the normal-ordered form, we investigate this approximation for the perturbative triples correction.
Results are compared with a code-generation based transcorrelated coupled cluster implementation up to quadruple excitations.
@@ 1 Introduction
The accurate description of electron correlation remains one of the greatest challenges in quantum chemistry.
Within a one-particle picture, the energy contribution of electron correlation can be defined as the difference between the exact full configuration interaction (FCI), and the mean-field Hartree–Fock (HF) energy.
One of the most powerful approaches to recover the correlation energy is the coupled clusterCoester_1958, Coester_1960, Cizek_1966, Cizek_1971 (CC) approach.
In CC, the wave function is parametrized by an exponential cluster operator acting on a reference wave function (usually the HF determinant).
In practice, the respective wave function expansion is truncated after single and double excitationsPurvis_1982, Scuseria_1987 (CCSD) to avoid the exponential scaling.
The inclusion of full triple excitations (CCSDT)Noga_1987, Scuseria_1988 can significantly improve on the accuracy of the correlation energyLesiuk_2022, but accounting for it is computationally too demanding for most system sizes.
Hence, triples are usually approximated by perturbative approachesRaghavachari_1989 (in particular in the CCSD(T) model) as a good compromise between cost and accuracy.
The size of the one-particle basis, given by a set of Gaussian functions, determines the overall quality of the a calculation.
With an increasing number of Gaussians, the basis set eventually converges to the complete basis set limit.
@@
However, due to the high computational scaling of most ab initio methods, it is not always feasible to apply large basis sets.
Explicitly correlated methods are a powerful alternative in this context as they enforce the electron-electron cusp condition Kato_1957, helgaker_molecular_2000 by explicitly incorporating the electronic distance into the wave function.
As a result, the convergence of the results toward the complete basis set limit is greatly accelerated.
For most applications, the ansätze of the R12/F12 family Haettig_2012, Kong_2012, Tenno_2012 have become standard tools of providing a computationally tractable way to introduce explicit correlation at the level of electronic pair functions.
However, a significant drawback of these methods is the nontrivial generalization to higher-than-double excitations.
This is an important deficiency as it is well-known that at least triple excitations are required to achieve chemical accuracy for most properties.
In fact, in most implementations of the F12 variant of the popular CCSD(T) method, triply excited contributions to the energy are evaluated using the standard orbital formula, i.e., without any form of explicit correlation.
In other words, only the CCSD-F12 component of the energy benefits from the accelerated convergence to the basis set limit, while the (T) component is not improved in a meaningful way.
Clearly, this is not a satisfactory situation as the basis set error in the (T) contribution will likely dominate in the overall error budget.
A natural route to solve this problem is to extend the F12 theory to triple (and possibly higher) excitations.
This approach was adopted by KöhnKohn2009, Koehn_2009_f12Triples, kohn2010 who proposed a modified ansatz for explicitly correlated CC wave functions which incorporates explicitly correlated components into the connected triple excitations using a set of cusp conditions (extended SP ansatz).
It was demonstrated that this method accelerates the convergence of the (T) correction to the complete basis set limit, but at a significantly increased cost in comparison with the standard approach.
A much simpler approach to address the aforementioned problem of lack of explicit correlation in the (T) correction is based on scaling it by the ratio of MP2-F12 and MP2 energies in a given basis (see Refs. Knizia_2009 and Adler_2007 and references therein for an extended discussion).
This procedure typically improves the results significantly, but is not free from drawbacks.
First, it breaks size consistency of the results which becomes problematic in applications to large systems.
Second, it assumes that the MP2 and (T) energies converge to the CBS limit at the same rate.
However, MP2 energies tend to converge more slowly, and hence, the scaled (T) correction is typically an overestimation.
These problems were addressed in the recent work of Kállay and collaborators kallay2021.
They propose to perform the scaling not at the level of the total (T) correction, but divide it into contributions from individual triplets of occupied orbitals () and scale each such contribution separately.
While this approach mitigates the size-consistency issue, it is still based on the assumption that MP2 and (T) energies converge to the CBS limit at the same rate.
A different approach to explicitly correlate electrons is transcorrelation, introduced by Boys and HandyBoys_Handy_1969_Transcorrelation, which we adopt in this study.
In contrast to including the interelectronic distance into the wave function, the Hamiltonian is similarity transformed by a correlation factor.
As a result, transcorrelation is generally applicable, even to multiconfigurational wave functions Baiardi_2020, Baiardi_2022, Cohen_2019, Liao_2023.
However, the resulting transcorrelated Hamiltonian is no longer Hermitian and includes a three-body operator.
Recently, the transcorrelation idea has experienced a vivid revivalNooijen_1998, Tenno_2000a, Tenno_2000b, Imamura_2003, Umezawa_2003, Umezawa_2004a, Umezawa_2004b, Umezawa_2005, Luo_2010a, Luo_2011, Luo_2012, Luo_2010b, Yanai_2012, Sharma_2014, Ochi_2012, Ochi_2014, Kersten_2016, Cohen_2019, Dobrautz_2019, Luo_2018, Gunter_2021, Ochi_2016, Ochi_2015, Jeszenszki_2018, Liao_2021, Schraivogel_2021, Schraivogel_2023, Luo_2022, Motta_2020, Kumar_2022, Khamoshi_2021, Giner_2021, Dobrautz_2022, Ammar_2022, Liao_2023, Ochi_2023, lee2023a, ammar2023, kats2024, szenes2024.
In the context of CC, the non-Hermiticity does not impose any issues, since the Hamiltonian is similarity transformed by the exponential cluster operator, resulting already in a non-Hermitian Hamiltonian.
Nevertheless, the full treatment of the three-body operator increases the computational cost and memory scaling.
Ten-no combined the transcorrelated Hamiltonian with linearized CCHino_2002, where many contractions involving the three-body operator were neglected.
Later Alavi and co-workers applied CCSD to the transcorrelated Hamiltonian (TC-CCSD) for the homogeneous electron gasLiao_2021, atomsSchraivogel_2021 and moleculesSchraivogel_2023.
We emphasize, however, that the basis-set limit of TC-CCSD is different from that of CCSD.
In order to deal with the three-body couplings, they introduced an approximation based on the normal-ordered HamiltonianSchraivogel_2023.
In this approximation, the full TC Hamiltonian is normal-ordered with respect to a single reference determinant and subsequently the remaining three-body integrals are neglected.
Within this procedure, mean-field contributions are transferred from the three-body operator to two-, one-, and zero-body operators.
In the same paper, they introduced a similar approximation based on the -dressed HamiltonianKoch_1994, DePrince_2013, which, however, still suffers from the huge memory scaling.
In the -dressed Hamiltonian approximation, all remaining three-body terms are neglected, meaning that effective intermediates are precontracted with the three-body operator before the operator is neglected.
Another approachChristlmaier_2023 to reduce the drawbacks of the three-body operator is based on the generalized normal-orderingkutzelnigg1997, kutzelnigg1999 of the three-body operator, which reduces the scaling to .
For higher-order CC models, the explicit treatment of the three-body operator results in numerous possible diagrams.
A derivation of the CCSD equation for Hamiltonians including three-body couplings was given by Piecuch and co-workers in the context of nuclear structure theoryHagen_2007.
The total number of CCSD diagrams for a Hamiltonian including three-body couplings increases from 48 to 116.
Hence, tools for automatic code contractions become important for the implementation of TC-CC with higher than doubles excitations.
@@
In this work, we derive a non-iterative triples approximation for TC-CCSD(T) including three-body operators and the normal-order approximation.
A first CCSD(T) implementation for Hamiltonians including a three-body operatorbinder2013 was derived based on the CCSD(T)kucharski1998, crawford1998, taube2008a method.
Later, Kats et al. derived the CCSD(T) method for the transcorrelated Hamiltoniankats2024, however without the inclusion of the full three-body operator.
We implemented TC-CCSDT and TC-CCSDTQ for the full transcorrelated Hamiltonian with the Wick&dEvangelista_2022, fevangelista_2023 library to generate the required CC tensor contractions.
In the following section, we first briefly review the conventional non-iterative triples approximation and the transcorrelated Hamiltonian.
Afterward, we show the required diagrams including three-body couplings for perturbative triples.
The correlation factor applied in this work is introduced in Section LABEL:Sec:Comp_Meth.
We compare these TC-CC models for Be, Ne, LiH and Be2 in Section LABEL:Sec:Results.
@@ 2 Theory
If not stated otherwise, we assume spin-orbital labels for all operators.
For convenience, we summarize our notation in Table 1.
symbol
meaning
occupied spin-orbital indices
virtual spin-orbital indices
general spin-orbital indices
denotes a normal-ordered string of second quantized operators
electron repulsion integrals in physics notation
antisymmetrized two-body integrals, e.g.
Table 1: Notation in the Current Work.
@@
is the mean-field energy contribution of the reference determinant, including the nuclear repulsion energy .
is an element of the Fock matrix
where comprises all one-body operators, i.e. the kinetic energy and electron-nuclear interactions.
In order to distinguish from the transcorrelated Hamiltonian, the former quantity is referred to as the conventional Hamiltonian.
We consistently map the indices of the creation and annihilation operators by placing them in the superscript and subscript, respectively, in the corresponding tensor (bra and ket of antisymmetrized integrals, respectively).
We employ the usual diagrammatic representationfeynman1949, goldstone1957, hugenholtz1957, where dashed lines represent operators.
For example, two outgoing arcs connected to an “x” correspond to a one-body operator, two outgoing arcs connected to another set of two outgoing arcs represent a two-body operator.
Solid horizontal lines represent amplitudes, where the number of vertices corresponds to the excitation degree of the amplitude.
The dotted line represents the denominator constructed from the diagonal entries of the Fock operator; for example
is the three-body denominator.
The CC amplitudes are denoted by the usual symbols , , , and so forth.
For convenience, we will employ both an algebraic and diagrammatic representation in this work.
@@ 2.1 Conventional Perturbative Triples
In this section, we rederive the conventional perturbative triples correction to the energy in the formalism of StantonStanton_1997.
In the subsequent section, this framework is extended to incorporate the three-body operator in the electronic Hamiltonian.
As described by Stanton in Ref. Stanton_1997, the CCSD method with can be transformed into an eigenvalue problem of , where the eigenvector is the reference determinant
which is valid after projection onto a manifold of singly- and doubly-excited determinants.
However, since the similarity transformation is not unitary, is not Hermitian.
Hence, the left eigenvector is not given by
In order to evaluate the left eigenvector explicitly, the so-called equations have to be solved, so that
with where is a de-excitation operator.
Now, instead of , is the zeroth-order Hamiltonian with and as left- and right-hand states, respectively.
The perturbative expansion of the similarity transformed Hamiltonian according to StantonStanton_1997 results in the leading-order correction to the energy
where S, D, and T correspond to singly, doubly, and triply excited determinants, respectively, and D3 is a diagonal matrix with the inverse of the elements defined in Eq. (4) on the diagonal.
The order of the Hamiltonian in the perturbation theory expansion is indicated by the square brackets, i.e. [1] or [2].
This perturbative expansion leads to two different sets of amplitudes which can be evaluated according to
@@
Note that due to the commutator in Eq. (11) only connected diagrams can be created, while in Eq. (12) also disconnected diagrams appear.
The tensor contractions for the first and second intermediate have the following form
@@
respectively.
Note that the last term in the latter equation vanishes due to the Brillouin condition in the case of conventional Hamiltonian, but this is no longer true for the transcorrelated Hamiltonian considered in the subsequent sections.
While the transition from Eq. (20) to Eq. (24) is in general not valid, we emphasize here that, due to the antisymmetry of the two-body integrals and the approximation of , Eqs. (20) and (24) are equivalent.
Therefore, in practice, the latter two diagrams in Eq. (20) are added to Eq. (15), so that the first tensor contraction must be evaluated only once.
The energy contribution of the non-iterative triples correction for the conventional Hamiltonian can be written as
The skeleton diagrams in Eq. (29) can be derived by contracting the intermediate amplitudes, and , with the denominator .
@@ 2.2 Transcorrelated Hamiltonian
The TC Hamiltonian can be derived by similarity transformation of the conventional Hamiltonian with a Jastrow factorjastrow1955 of the form
resulting in the additional two-body term and the three-body term .
For the sake of completeness, we summarize some of the important aspects of the TC Hamiltonian.
In the following, we redefine